# Learning to Explore with Lagrangians for Bandits under Unknown Linear Constraints

### v1

@@ Abstract
Pure exploration in bandits models multiple real-world problems, such as tuning hyper-parameters or conducting user studies, where different safety, resource, and fairness constraints on the decision space naturally appear. We study these problems as pure exploration in multi-armed bandits with unknown linear constraints, where the aim is to identify an $r$-good feasible policy. First, we propose a Lagrangian relaxation of the sample complexity lower bound for pure exploration under constraints. We show how this lower bound evolves with the sequential estimation of constraints. Second, we leverage the Lagrangian lower bound and the properties of convex optimisation to propose two computationally efficient extensions of Track-and-Stop and Gamified Explorer, namely LATS and LAGEX. To this end, we propose a constraint-adaptive stopping rule, and while tracking the lower bound, use pessimistic estimate of the feasible set at each step. We show that these algorithms achieve asymptotically optimal sample complexity upper bounds up to constraint-dependent constants. Finally, we conduct numerical experiments with different reward distributions and constraints that validate efficient performance of LAGEX and LATS with respect to baselines.
@@
1 INTRODUCTION
Decision-making under uncertainty is a ubiquitous challenge encountered across various domains, including clinical trials (Villar et al.,, 2015), recommendation systems (Zhao and Yang,, 2024), and more. Multi-Armed Bandit (MAB) serves as an archetypal framework for sequential decision-making under uncertainty and allows us to study the involved information-utility trade-offs (Lattimore and Szepesvári,, 2020). In MAB, at each step, an agent interacts with an environment consisting of $K$ decisions (also knows as arms) corresponding to $K$ noisy feedback distributions (or reward distributions). At each step, the agent takes a decision, and obtains a reward from its unknown reward distribution. The goal of the agent is to compute a policy, i.e. a distribution over the decisions, that maximises a certain utility metric (e.g. accumulated rewards (Auer et al.,, 2002), probability of identifying the best arm (Kaufmann et al.,, 2016) etc.).
In this paper, we focus on the pure exploration problem of MABs, where the agent interacts by realising a sequence of policies (or experiments) with the goal of answering a query as correctly as possible. A well-studied pure exploration problem is Best-Arm Identification (BAI), where the agent aims to identify the arm with the highest expected reward (Even-Dar et al., 2002a, ; Bubeck et al.,, 2009; Jamieson and Nowak,, 2014; Kaufmann et al.,, 2016). BAI has been applied in hyper-parameter tuning (Li et al.,, 2017), communication networks (Lindståhl et al.,, 2022), influenza mitigation (Libin et al.,, 2019), finding the optimal dose of a drug (Aziz et al., 2021a, ) etc. However, real-world scenarios often impose constraints on the arms that must be satisfied (Carlsson et al.,, 2024). For example, Baudry et al., (2024) considers a recommendation problem with the aim to guarantee a fixed (known) minimum expected revenue per recommended content while identifying the best content from bandit feedback. Additionally, if we have multiple objectives in a decision making problem, a popular approach to optimize them is finding the optimal policy for one objective while constraining the others (Fonseca and Fleming,, 1998). In the spirit of Example 1 of (Carlsson et al.,, 2024), let us consider another example of a person searching for the most tasty meal plan given a bunch of available food items. The person’s aim is also to gain enough nutrition by eating a certain amount of carbohydrate, protein, fat, and fibres. This problem can be modelled as a pure exploration problem, where one tries different combinations and fractions of food items to find the most tasty combination while restricting the nutritional intakes over a recommended threshold.
Pure Exploration under Constraints. In recent years, such aforementioned real-life problems naturally motivated the study of pure exploration under a set of known and unknown constraints (Katz-Samuels and Scott,, 2018; Wang et al., 2021b, ; Li et al.,, 2023; Wu et al.,, 2023; Carlsson et al.,, 2024). Specifically, we aim to find the optimal policy that maximises the expected rewards over the set of arms and also satisfies the true constraints with confidence $1-\delta$. This is known as the fixed-confidence setting of pure exploration (Wang et al., 2021b, ; Carlsson et al.,, 2024), while there exists fixed-budget setting also which is of independent interest (Katz-Samuels and Scott,, 2018; Li et al.,, 2023; Faizal and Nair,, 2022). Existing literature has studied either the general linear constraints when they are known (Carlsson et al.,, 2024; Camilleri et al., 2022a, ), or very specific type of unknown constraints, e.g. safety (Wang et al., 2021b, ), knapsack (Li et al.,, 2023), fairness (Wu et al.,, 2023), preferences (Lindner et al.,, 2022) etc. Here, we study the pure exploration problem in the fixed-confidence setting subject to unknown linear constraints on the policy, which generalises all these settings (Section 2). Further discussions on related works is in Appendix B.1.
Recently, Carlsson et al., (2024) shows that if the constraints are known, and so is the feasible policy set, a bandit instance may become harder or easier depending on the geometry of constraints. They state that studying similar phenomenon for unknown constraints is an open problem as the feasible policy set has to be estimated. Specifically, we have to simultaneously control concentrations of the mean rewards and the constraints to their ‘true’ values to reach the feasible policy set. Additionally, at any finite time, the estimated constraints might exhibit small but non-zero errors. This prevents from ‘exactly’ obtaining the feasible policy set, and consequently, the recommended optimal and feasible policy (though the error might be below numerical limits). Hence, for rigour, we relax the problem of finding the optimal feasible policy to finding an $r$-good feasible policy. For a given $r>0$, an $r$-good policy has mean reward not more than $r$ away from that of the optimal policy (Mason et al.,, 2020; Jourdan and Degenne,, 2022; Jourdan et al.,, 2023). In practice, one can consider $r$ to be the numerical limit of computation or a reasonably small quantity.
This leads us to two questions
  • How does the hardness of pure exploration to find $r$-good feasible policy under unknown constraints change if the constraints are estimated sequentially?
  • How can we design a generic algorithmic scheme to track both the constraints and optimal policy with sample- and computational-efficiency?
Our Contributions address these questions as follows:
1. Lagrangian relaxation of the lower bound. The minimum number of samples required to conduct $r$-good pure exploration with fixed confidence is expressed by a lower bound, which is an optimisation problem under known constraints. To efficiently handle unknown constraints, we propose a novel Lagrangian relaxation of this optimisation problem (Section 3). At every step, we construct pessimistic feasible policy set, and plug it in the lower bound. The Lagrangian multipliers balance the identifiability of an $r$-good policy and the feasibility under estimated constraints. We leverage results from convex analysis to show that the Lagrangian relaxation of the lower bound with pessimistic feasible set preserves all the desired properties of the lower bound under known constraints, and consequently, allow us to design lower bound tracking algorithms. Further, we characterise the Lagrangian lower bound for Gaussian rewards, which connects with the lower bound for known constraints.
2. A generic algorithm design. We leverage this Lagrangian formulation of lower bound to propose two algorithms. First, incorporating the constraint estimates in the existing stopping rule for known constraints ensures concentration of both the means’ and constraints’ estimates to their true values before recommending an $r$-good policy with confidence level $1-\delta$. We show that this stopping rule also ensures that the recommended policy is feasible with the same confidence level. Then, we extend the Track-and-Stop (Garivier and Kaufmann,, 2016) and gamified explorer (Degenne et al., 2019b, ) approaches for the Lagrangian lower bound to design LATS (Lagrangian Track and Stop) and LAGEX (Lagrangian Gamified EXplorer), respectively (Section 4).
3. Upper bound on sample complexities. We provide upper bounds on sample complexities of LATS and LAGEX (Section 4). This requires proving a novel concentration inequality for the constraints. As a consequence, LATS achieves an upper bound, which is $(1+\mathfrak{s})$ times the upper bound of TS under known constraints. $\mathfrak{s}$ is the shadow price of the true constraints and quantifies its stability under perturbation. In contrast, LAGEX leads to an upper bound that has only an additive $\mathfrak{s}$ factor with the known constraint lower bound. This suggests that LAGEX should be more sample-efficient than LATS. Our experimental results across synthetic and real data validate that LAGEX requires the least samples among competing algorithms and exactly tracks the change in hardness due to constraints across environments (Section 5).
2 PROBLEM FORMULATION
Notation. $x$, $\mathbf{x}$, $\mathbf{X}$, and $\mathcal{X}$ denote a scalar, a vector, a matrix, and a set respectively. $\mathbf{x}_{i}$ denotes $i$-th component of $\mathbf{x}$. For a positive semi-definite matrix $\mathbf{A}$ and vector $\mathbf{z}$, $\|\mathbf{z}\|_{\mathbf{A}}^{2}=\langle\mathbf{z},\mathbf{A}\mathbf{z}\rangle$. Also, in $\mathbb{R}_{+}^{d}$, we include $\mathbf{0}^{d}$. $[K]$ refers to $\{1,\ldots,K\}$. $\mathrm{Supp}(P)$ denotes the support of a distribution $P$. $\Delta_{K}$ is the simplex over $[K]$.
Problem Formulation. We work with a MAB instance with $K\in\mathbb{N}$ arms. Each arm $a\in[K]$ has a reward distribution $P_{a}$ with unknown mean ${\boldsymbol{\mu}}_{a}\in\mathbb{R}$. The agent, at each time step $t\in\mathbb{N}$, chooses an action $A_{t}\in[K]$, and observes a stochastic reward $R_{t}\sim P_{A_{t}}$. A feasible policy $\boldsymbol{\pi}\in\Delta_{K}$ satisfies $\mathbf{A}\boldsymbol{\pi}\leq\mathbf{0}$ with respect to the set of $d$ linear constraints $\mathbf{A}\in\mathbb{R}^{d\times{K}}$. [^1] We assume that simplex constraints are augmented in $\mathbf{A}$ for ease of notation.
If $\mathbf{A}$ is known, the agent has access to the non-empty and compact set of feasible policies $\mathcal{F}\mathrel{\triangleq}\{\boldsymbol{\pi}\in\Delta_{K}\mid\mathbf{A}\boldsymbol{\pi}\leq\boldsymbol{0}\}$. The agent aims to identify an $r$-good optimal feasible policy, i.e. any feasible policy which belongs to the set $\Pi_{\mathcal{F}}^{r}\mathrel{\triangleq}\{\boldsymbol{\pi}\in\mathcal{F}\mid{\boldsymbol{\mu}}^{T}\boldsymbol{\pi}+r\geq{\boldsymbol{\mu}}^{T}\boldsymbol{\pi}_{\mathcal{F}}^{\star}\}$, where
$\boldsymbol{\pi}^{\star}_{\mathcal{F}}\mathrel{\triangleq}\operatorname*{arg\,max}_{\boldsymbol{\pi}\in\mathcal{F}}{\boldsymbol{\mu}}^{T}\boldsymbol{\pi}\,.$ (1)
Definition 1 ($(1-\delta)$-correct and $(1-\delta)$-feasible $r$-good pure exploration).
For $\delta\in[0,1)$, a policy recommended by a pure exploration algorithm is $(1-\delta)$-correct and $(1-\delta)$-feasible if $\Pr[\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}\neq\boldsymbol{\pi}^{\star}_{\mathcal{F}}]\leq\delta$ and $\Pr[{\mathbf{A}}\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}\geq\boldsymbol{0}]\leq\delta$.
In our setting, we do not have access to the true set of constraints. Hence, using the observations, we construct $\hat{\mathbf{A}}$ as an estimate of $\mathbf{A}$. Then, the agent builds an estimated feasible set ^$\hat{\mathcal{F}}\mathrel{\triangleq}\{\boldsymbol{\pi}\in\Delta_{K}\mid\hat{\mathbf{A}}\boldsymbol{\pi}\leq\textbf{0}\}$ to identify the optimal feasible policy as $\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}\mathrel{\triangleq}\arg\max_{\boldsymbol{\pi}\in\hat{\mathcal{F}}}{\boldsymbol{\mu}}^{T}\boldsymbol{\pi}$. In addition, the estimated $r$-good policy set is $\Pi_{\hat{\mathcal{F}}}^{r}\mathrel{\triangleq}\{\boldsymbol{\pi}\in\hat{\mathcal{F}}\mid{\boldsymbol{\mu}}^{T}\boldsymbol{\pi}+r\geq{\boldsymbol{\mu}}^{T}\boldsymbol{\pi}_{\hat{\mathcal{F}}}^{\star}\}$. We know that obtaining accurate estimates of these quantities would require us to collect feedback of satisfying constraints over time. This poses an additional cost on top of using observations to estimate ${\boldsymbol{\mu}}$.
@@ 2 PROBLEM FORMULATION
Goal. In order to recommend a $(1-\delta)$-correct and $(1-\delta)$-feasible policy that is $r$-good with respect ot the true optimal policy $\boldsymbol{\pi}^{\star}_{\mathcal{F}}$, we aim to minimise the expected number of interactions $\mathbb{E}[\tau_{\delta}]\in\mathbb{N}$.
@@
Motivation: Extension of Prior Setups. Now, we clarify our motivation by showing that different existing problems are special cases of our setting.
@@
a. Thresholding Bandits. Thresholding bandits (Aziz et al., 2021a, ) are motivated from the safe dose finding problem, where one wants to identify the most effective dose of a drug below a known safety level. This has also motivated the safe arm identification problem (Wang et al., 2021b, ). Our setting generalises it further and detects the optimal combination of doses of available drugs yielding highest efficacy while staying below the safety threshold. Formally, we identify $\boldsymbol{\pi}^{\star}=\operatorname*{arg\,max}_{\boldsymbol{\pi}\in\Delta_{K}}{\boldsymbol{\mu}}^{T}\boldsymbol{\pi}$, such that $\mathbf{I}\boldsymbol{\pi}\leq\mathbf{I}{\boldsymbol{\theta}}$ for thresholds ${\boldsymbol{\theta}}_{a}>0$.
b. Optimal Policy under Knapsack. Bandits under knapsack constraints are studied in both BAI (Li et al.,, 2023; Tran-Thanh et al.,, 2012; Li et al.,, 2021) and regret minimisation literature (Badanidiyuru et al.,, 2018; Agrawal and Devanur,, 2016; Immorlica et al.,, 2022; Agrawal et al.,, 2016; Sankararaman and Slivkins,, 2018). Detecting an optimal arm might have additional resource constraints than the number of required samples. This led to study of BAI with knapsacks under fixed-budget setting (Li et al.,, 2023). But one might only want to deploy a final policy that maximises utility and satisfies knapsack constraints. For example, we want to manage caches where the recommended memory allocation should satisfy a certain resource budget but can violate them during exploration (End-of-Time constraint (Carlsson et al.,, 2024)). Formally, $\boldsymbol{\pi}^{\star}_{\tau_{\delta}}=\operatorname*{arg\,max}_{\boldsymbol{\pi}\in C_{\mathbf{A}}}\hat{{\boldsymbol{\mu}}}_{\tau_{\delta}}^{T}\boldsymbol{\pi}$, and $C_{\mathbf{A}}\mathrel{\triangleq}\{\mathbf{A}\boldsymbol{\pi}_{\tau_{\delta}}\leq c,c\geq 0\}$.
@@
c. BAI with Fairness Constraints across Sub-populations (BAICS). Wu et al., (2023) aim to select an arm that must be fair across $M$ sub-populations rather than the whole population as in standard BAI. Finding only the optimal arm might not be enough because it might not perform equally good for all of the sub-populations. Here, the arm belongs to a set $C:=\{k\in[{K}]|{\boldsymbol{\mu}}_{k,m}\geq 0,m\in[M]\}$ where the observation for arm $k$ and population $m$ comes from $\mathcal{N}(\mu_{k,m},1)$. It ensures that the chosen arm does not perform too bad for any sub-population. This is similar to the problem where we have $M$ groups of patients and ${K}$ drugs to administer with reward means ${\boldsymbol{\mu}}_{k}$. Then we are looking for a mixture of drugs to administer such that $\boldsymbol{\pi}^{\star}=\operatorname*{arg\,max}_{\boldsymbol{\pi}\in\Delta_{K}}{\boldsymbol{\mu}}^{T}\pi$, such that $\mathds{1}_{{\boldsymbol{\mu}}_{m}\geq\boldsymbol{0}}^{\top}\boldsymbol{\pi}=1,\forall m\in[M]$. Hence, our setting models BAICS as a special case.

### v2

# Learning to Explore with Lagrangians for Bandits under Unknown Constraints

@@ Abstract
Pure exploration in bandits formalises multiple real-world problems, such as tuning hyper-parameters or conducting user studies to test a set of items, where different safety, resource, and fairness constraints on the decision space naturally appear. We study these problems as pure exploration in multi-armed bandits with unknown linear constraints, where the aim is to identify an $r$-optimal and feasible policy as fast as possible with a given level of confidence. First, we propose a Lagrangian relaxation of the sample complexity lower bound for pure exploration under constraints. Second, we leverage properties of convex optimisation in the Lagrangian lower bound to propose two computationally efficient extensions of Track-and-Stop and Gamified Explorer, namely LATS and LAGEX. Then, we propose a constraint-adaptive stopping rule, and while tracking the lower bound, use optimistic estimate of the feasible set at each step. We show that LAGEX achieves asymptotically optimal sample complexity upper bound, while LATS shows asymptotic optimality up to novel constraint-dependent constants. Finally, we conduct numerical experiments with different reward distributions and constraints that validate efficient performance of LATS and LAGEX.
1 Introduction
Multi-Armed Bandit (MAB) serves as an archetypal framework for sequential decision-making under uncertainty that allows us to study the corresponding information-utility trade-offs (Lattimore and Szepesvári,, 2020). In MAB, at each step, an agent interacts with an environment consisting of $K$ decisions (also knows as arms) corresponding to $K$ noisy feedback distributions (or reward distributions). At each step, the agent takes a decision, and obtains a reward from its unknown reward distribution. The goal of the agent is to compute a policy, i.e. a distribution over the decisions, maximising a utility metric (e.g. accumulated rewards (Auer et al.,, 2002), probability of identifying the best arm (Kaufmann et al.,, 2016) etc.).
In this paper, we focus on the pure exploration problem of MABs, where the agent interacts by realising a sequence of policies (or experiments) with the goal of answering a query about the environment as correctly as possible. A well-studied pure exploration problem is Best-Arm Identification (BAI), where the agent aims to identify the arm with highest expected reward (Even-Dar et al., 2002a, ; Bubeck et al.,, 2009; Jamieson and Nowak,, 2014; Kaufmann et al.,, 2016). Kaufmann et al., (2016) derives an information-theoretic lower bound quantifying the minimum number of agent-environment interactions needed to identify the best arm with a given level of confidence. The lower bound depends on optimising the weighted sum of KL-divergences between the reward distributions of arms and their most confusing counterparts (Equation (3)). In this paper, we leverage the lower bound for algorithm design that plugs-in the empirical estimates of the reward distributions in lower bound and solves the optimisation problem on-the-go (Kaufmann et al.,, 2016; Degenne et al., 2019b, ; Carlsson et al.,, 2024).
BAI has been applied in hyper-parameter tuning (Li et al.,, 2017), communication networks (Lindståhl et al.,, 2022), influenza mitigation (Libin et al.,, 2019), finding the optimal dose of a drug (Aziz et al., 2021a, ) etc. However, real-world scenarios often impose constraints on the arm pulls (Carlsson et al.,, 2024). For example, Baudry et al., (2024) considers a recommendation problem with the aim to guarantee a fixed (known) minimum expected revenue per recommended content while identifying the best content from bandit feedback. Additionally, if we have multiple objectives in a decision making problem, a popular approach to optimize them is finding the optimal policy for one objective while constraining the others (Fonseca and Fleming,, 1998).
Pure Exploration under Constraints. The aforementioned problems motivated the study of pure exploration under a set of known and unknown constraints (Katz-Samuels and Scott,, 2018; Wang et al., 2021b, ; Li et al.,, 2023; Wu et al.,, 2023; Carlsson et al.,, 2024). Specifically, we aim to find the optimal policy that maximises the expected reward obtained from the set of arms and satisfies the true constraints, with confidence $1-\delta$. This is known as the fixed-confidence setting of pure exploration (Wang et al., 2021b, ; Carlsson et al.,, 2024). On the other hand, there is also the fixed-budget setting, which is of independent interest (Katz-Samuels and Scott,, 2018; Li et al.,, 2023; Faizal and Nair,, 2022). Existing literature has studied either the general linear constraints when they are known (Carlsson et al.,, 2024; Camilleri et al., 2022a, ), or very specific type of unknown constraints, e.g. safety (Wang et al., 2021b, ), knapsack (Li et al.,, 2023), fairness (Wu et al.,, 2023), preferences (Lindner et al.,, 2022) etc. Here, we study the pure exploration problem in the fixed-confidence setting subject to unknown linear constraints on the policy, which generalises all these settings (Section 2). A detailed discussion on related works is deferred to Appendix B.1.
Recently, Carlsson et al., (2024) derives a tight lower bound and designs asymptotically optimal algorithms for this problem when the constraints are known, and show that a bandit instance might become harder or easier depending on the geometry of the constraints. They pose that studying similar phenomenon for unknown constraints is an open problem as constraints are also estimated.
The challenge is that the lower bound dictating the hardness of the constrained pure exploration problem is sensitive to the active constraints, and for unknown constraints, we only have access to estimated constraints with non-zero error. This affects the exact detection of the optimal feasible policy.
Example. Let us consider a bandit environment with 4 arms with Gaussian rewards – means ${\boldsymbol{\mu}}=[1.0,0.8,0.6]$ and variance $1$. The constraints are $A{\boldsymbol{\pi}}\leq\mathbf{0}$ where $A=\begin{bmatrix}0.0,-0.3,-0.2\\ -0.2,0.1,0.1\end{bmatrix}$. Here, the optimal policy is ${\boldsymbol{\pi}}^{\star}=[1.0,0,0]$, and the first constraint is active at the optimal policy. Figure 1 shows a perturbation of the active constraint by $0.005$ shifts the optimal policy, and the sample complexity (Carlsson et al.,, 2024) blows with $\mathcal{O}(4\times 10^{4})$.
Figure 1: Effect of rank-one update of active constraint
Hence, for rigour, we relax the problem of finding the optimal feasible policy to finding an $r$-optimal feasible policy. For a given $r>0$, an $r$-optimal policy has mean reward not more than $r$ away from that of the optimal policy (Mason et al.,, 2020; Jourdan and Degenne,, 2022; Jourdan et al.,, 2023). This leads us to two questions:
1. How does the hardness of finding $r$-optimal feasible policy change under unknown constraints?
2. How can we design a generic algorithmic scheme to track both the constraints and optimal policy with sample- and computational-efficiency?
Our Contributions positively address the questions.
1. Lagrangian relaxation of the lower bound. Minimum number of samples required to conduct $r$-optimal pure exploration with fixed confidence is expressed by a lower bound– an optimisation problem under known constraints (Eq. (3)). To efficiently handle unknown constraints, we propose a novel Lagrangian relaxation of this optimisation (Section 3). At every step, we construct optimistic feasible policy set, and plug it in the relaxation. Lagrangian multipliers balance the identifiability of an $r$-optimal policy and the feasibility under estimated constraints. We leverage results from convex analysis to show that the relaxed lower bound with optimistic feasible set preserves all the desired properties of the lower bound under known constraints, and thus, allows designing lower bound tracking algorithm.
2. Generic algorithm design. First, we propose a new stopping rule accommodating the estimated constraints.This ensures concentration of the estimates of mean rewards and constraints to their true values before final recommendation. We further show this stopping rule recommends a policy that is both feasible and $r$-optimal with confidence at least $1-\delta$. Then, we extend the Track-and-Stop (Garivier and Kaufmann,, 2016) and gamified explorer (Degenne et al., 2019b, ) approaches with the Lagrangian lower bound to design LATS (LAgrangian Track and Stop) and LAGEX (LAgrangian Gamified EXplorer), respectively in (Section 4).
3. Upper bounds on sample complexities. We derive upper bounds on the sample complexities of LATS and LAGEX (Section 4). This requires proving a novel concentration inequality for the constraint estimates. As a consequence, LATS achieves an upper bound, which is $(1+\mathfrak{s})$ times the asymptotic upper bound of TS under known constraints, while LAGEX exhibits optimality. $\mathfrak{s}$ is the shadow price (ratio between maximum and minimum index value of the slack vector) that quantifies its stability under perturbation. Finally, we conduct experiments across synthetic and real data. We observe that LAGEX requires the least samples among competing algorithms and exactly tracks the change in hardness due to constraints across environments (Section 5).
2 Exploration under Unknown Constraints
Problem Formulation. We work with a MAB instance consisting of $K\in\mathbb{N}$ arms. Each arm $a\in[K]$ has a reward distribution $P_{a}$ with unknown mean ${\boldsymbol{\mu}}_{a}\in\mathbb{R}$. At each step $t\in\mathbb{N}$, the agent chooses an action $A_{t}\in[K]$, and observes a stochastic reward $R_{t}\sim P_{A_{t}}$. A feasible policy $\boldsymbol{\pi}\in\Delta_{K}$ satisfies $A\boldsymbol{\pi}\leq\mathbf{0}$ with respect to $d$ linear constraints $A\in\mathbb{R}^{d\times{K}}$. [^1] Notations: $\mathbf{x}$, $X$, and $\mathcal{X}$ denote a vector, a matrix, and a set, respectively. Detailed notations are in Table 1. We augment the simplex constraints in $A$, and normalise each row of $A$, i.e. $\|A_{i}\|_{2}=1$ for all $i\in[d]$.
Algorithm 1 Pure Exploration in Bandits with Unknown Linear Constraints
1:Input: Tolerance $r>0$, Confidence level $\delta\in(0,1)$
2:for $t=1,\,\ldots$ do
3:  Decision/sampling: Play an arm $a_{t}\in[1,K]$
4:  Reward-cost Feedback: Observe reward $r_{t}\sim P_{a_{t}}$ and cost $\mathbf{c}_{t}\sim Aa_{t}+\eta_{t}$
5:  if more than $1-\delta$ confident about the estimated answer being correct then
6:   Stop and stopping time $\tau_{\delta}\leftarrow t$
7:  end if
8:end for
9:Recommendation: $\hat{\boldsymbol{\pi}}=\operatorname*{arg\,max}_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}_{\tau_{\delta}}}^{r}}\hat{{\boldsymbol{\mu}}}_{\tau_{\delta}}^{\top}\boldsymbol{\pi}$
If $A$ is known, the agent has access to the non-empty and compact set of feasible policies $\mathcal{F}\mathrel{\triangleq}\{\boldsymbol{\pi}\in\Delta_{K}\mid A\boldsymbol{\pi}\leq\boldsymbol{0}\}$. The agent aims to identify an $r$-optimal optimal feasible policy, i.e. any feasible policy which belongs to $\Pi_{\mathcal{F}}^{r}\mathrel{\triangleq}\{\boldsymbol{\pi}\in\mathcal{F}\mid{\boldsymbol{\mu}}^{\top}\boldsymbol{\pi}+r\geq{\boldsymbol{\mu}}^{\top}\boldsymbol{\pi}_{\mathcal{F}}^{\star}\}$, given
$\boldsymbol{\pi}^{\star}_{\mathcal{F}}\mathrel{\triangleq}\operatorname*{arg\,max}_{\boldsymbol{\pi}\in\mathcal{F}}{\boldsymbol{\mu}}^{\top}\boldsymbol{\pi}\,.$ (1)
Definition 1 ($(1-\delta)$-correct and $(1-\delta)$-feasible $r$-optimal pure exploration).
For $\delta\in[0,1)$, an $r$-optimal pure exploration algorithm is called $(1-\delta)$-correct and $(1-\delta)$-feasible if the policy $\hat{\boldsymbol{\pi}}$ recommended by it satisfies $\Pr[\hat{\boldsymbol{\pi}}\notin\Pi_{\mathcal{F}}^{r}]\leq\delta$ and $\Pr[{A}\hat{\boldsymbol{\pi}}\geq\boldsymbol{0}]\leq\delta$.
@@ 2 Exploration under Unknown Constraints
Goal. In order to recommend a $(1-\delta)$-correct and $(1-\delta)$-feasible policy that is $r$-optimal with respect ot the true optimal policy $\boldsymbol{\pi}^{\star}_{\mathcal{F}}$, we aim to minimise the expected number of interactions $\mathbb{E}[\tau_{\delta}]\in\mathbb{N}$.
2.1 Extension of Prior Bandit Problems
@@
Now, we clarify our motivation by showing that different existing problems are special cases of our setting.
@@
a. Thresholding Bandits. Thresholding bandits (Aziz et al., 2021a, ) are motivated from the safe dose finding problem, where one wants to identify the most effective dose of a drug below a known safety level. This has also motivated the safe arm identification problem (Wang et al., 2021b, ). Our setting generalises it further to detect the optimal combination of doses of available drugs yielding highest efficacy while staying below the safety threshold. Formally, we identify $\boldsymbol{\pi}^{\star}=\operatorname*{arg\,max}_{\boldsymbol{\pi}\in\Delta_{K}}{\boldsymbol{\mu}}^{\top}\boldsymbol{\pi}$, such that $I\boldsymbol{\pi}\leq I{\boldsymbol{\theta}}$ for thresholds ${\boldsymbol{\theta}}_{a}>0$.
@@
b. BAI with Fairness Constraints across Sub-populations (BAICS). Wu et al., (2023) aims to select an arm that must be fair across $M$ sub-populations. Here, the arm belongs to a set $C\triangleq\{k\in[{K}]|{\boldsymbol{\mu}}_{k,m}\geq 0,m\in[M]\}$ where the observation for arm $k$ and population $m$ comes from $\mathcal{N}({\boldsymbol{\mu}}_{k,m},1)$. It ensures that the chosen arm does not perform too bad for any sub-population. Like standard BAI, finding only the optimal arm might not be enough because it might not perform equally good for all of the sub-populations. This is similar to having $M$ groups of patients and K{K} drugs to administer, where we are looking for a mixture of drugs such that $\boldsymbol{\pi}^{\star}=\operatorname*{arg\,max}_{\boldsymbol{\pi}\in\Delta_{K}}{\boldsymbol{\mu}}^{\top}\boldsymbol{\pi}$, such that $\mathds{1}_{{\boldsymbol{\mu}}_{m}\geq\boldsymbol{0}}^{\top}\boldsymbol{\pi}=1,\forall m\in[M]$.
We refer to Section B.2 for further discussion on generalisation to the existing bandit settings.